OpenStudy (anonymous):

HELP in 6th grade Identify the property shown. 3p · (2 + k) = 6p + 3pk A. transitive B. associative C. commutative D. distributive

OpenStudy (anonymous):

It's distributive D.

OpenStudy (akashdeepdeb):

Take an example a(b+c) = ab + ac Which is the distributive property of arithmetic.

OpenStudy (anonymous):

Can you help me with 4 more questions

OpenStudy (akashdeepdeb):

a + (b+c) = (a+b) + c is the associative property of arithmetic. a + b = b + a is the commutative property.

OpenStudy (akashdeepdeb):

Have you tried them yourself?

OpenStudy (anonymous):

yes and i flunked

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

Use properties to evaluate the expression. (19 + 28) + 21 A. 48 B. 58 C. 68 D. 89

OpenStudy (anonymous):

i think it is C am i right

OpenStudy (akashdeepdeb):

yes

OpenStudy (anonymous):

so am i right? thanks here is another

OpenStudy (anonymous):

Evaluate the expression. 17 + (29 + 13) A. 49 B. 59 C. 69 D. 76

OpenStudy (anonymous):

i think it is B and am i Right?

OpenStudy (akashdeepdeb):

yes. :D

OpenStudy (anonymous):

here is another

OpenStudy (anonymous):

Which statement shows the distributive property? A. 13 · (k + 10) = 13 · k + 13 · 10 B. 13 · (k + 10) = 13 · k + 10 C. 13 · (k + 10) = 13 · k · 10 D. 13 · (k + 10) = (k + 10) · 13

OpenStudy (anonymous):

i have no answer i failed with this question

OpenStudy (anonymous):

do you know this one?

OpenStudy (akashdeepdeb):

Read the explanation for the distributive property I gave above.

OpenStudy (akashdeepdeb):

Say you have 10 (10 + 15) Now what this actually means is that: 10 (10 + 15) = 10*10 + 10*15 = 100 + 150 = 250 Or 10 (10 + 15) = 10 * 25 = 250 Which is the same. Thus it is just a property in Arithmetic. Which one do YOU think is the answer? :D

OpenStudy (anonymous):

A

OpenStudy (akashdeepdeb):

good. :)

OpenStudy (anonymous):

@AkashdeepDeb I have to admit.. you explain things well lol

OpenStudy (akashdeepdeb):

lol Thanks!

OpenStudy (anonymous):

it is right ya-ya one last question

OpenStudy (anonymous):

If m = s and s = w, which statement is true? A. m > w B. m = w C. m < w D. m + s = w

OpenStudy (anonymous):

i think it is B

OpenStudy (akashdeepdeb):

Now let us say the: No. of candies I have = 10 No. of candies = No. of candies you have. Which of the following is true? # No. of candies you have = 10 # No. of candies you have > No. of candies I have # No. of candies you have < No. of candies I have # No. of candies you have = 0.

OpenStudy (anonymous):

is B right /?

OpenStudy (akashdeepdeb):

What do you think is the answer to my question?

OpenStudy (anonymous):

oh sorry okay ummm the second one

OpenStudy (anonymous):

i feel like a dummy because i don't know the answer

OpenStudy (akashdeepdeb):

It is a pretty straight forward question. See do NOT take math like something that consists of A,B,C,D options. Math is in everyday life. So take an example: No. of candies I have = 10 No. of candies I have = No. of candies you have. Now how many candies do you have?

OpenStudy (akashdeepdeb):

well, the answer to your question IS B.

OpenStudy (anonymous):

i passed my test A +

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

here is your medal